### v1

@@ 6 Ordinals of size $\omega_{1}$
If ${\mathfrak{h}}=\omega_{1}<\omega_{2}=2^{\omega}=2^{\omega_{1}}$ and $|\delta|=\omega_{1}$, then
${\mathbb{P}}(\omega^{\delta})\equiv\left\{\begin{array}[]{cc}\mathop{\rm Col}\nolimits(\omega_{1},\omega_{2})\equiv(P(\omega)/\mathop{\rm Fin}\nolimits)^{+},&\mbox{ if (A) or (B);}\\[2.84526pt] \mathop{\rm Col}\nolimits(\omega,\omega_{2}),&\mbox{ if (D) or (E).}\end{array}\right.$
By Lemma 4.1 we have the following cases ((C) is impossible):
Case (A): $\mathop{\rm cf}\nolimits\delta\leq\omega$. Then by Theorem 4.10 and Fact 2.3 we have ${\mathbb{P}}(\omega^{\delta})\equiv\langle[\omega]^{\omega},\subset^{*}\rangle\ast\pi$. So, by Fact 2.8(e), ${\mathbb{P}}(\omega^{\delta})$ and, hence, $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ collapses $\omega_{2}$ to $\omega_{1}$. By Theorem 4.10 $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\omega_{1}$-closed and, by our assumption, $\omega_{2}^{<\omega_{1}}=\omega_{2}$ so, by Fact 2.6, ${\mathbb{P}}(\omega^{\delta})\equiv\mathop{\rm Col}\nolimits(\omega_{1},\omega_{2})\equiv(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$.
Case (B): $\delta=\theta+\omega_{1}$, where $\mathop{\rm cf}\nolimits\theta=\omega<\omega_{1}<\theta<\omega_{2}$. Then by Theorem 4.10 $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ is $\omega_{1}$-closed and ${\mathbb{P}}(\omega^{\delta})\equiv\langle[\omega]^{\omega},\subset^{*}\rangle\ast\pi$. So, as in Case (A) we have ${\mathbb{P}}(\omega^{\delta})\equiv\mathop{\rm Col}\nolimits(\omega_{1},\omega_{2})\equiv(P(\omega)/\mathop{\rm Fin}\nolimits)^{+}$.
Case (D) or (E). Then from Theorem 4.10 and Fact 2.3 it follows that ${\mathbb{P}}(\omega^{\delta})\equiv\langle[\omega_{1}]^{\omega_{1}},\subset^{*}\rangle\ast\pi$. So, by Fact 2.8(c), ${\mathbb{P}}(\omega^{\delta})$ and, hence, $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\delta})$ collapses $\omega_{2}$ to $\omega$. By Fact 2.6 we have ${\mathbb{P}}(\omega^{\delta})\equiv\mathop{\rm Col}\nolimits(\omega,\omega_{2})$.
Acknowledgement.

### v2

@@ 5 The posets ${\mathbb{P}}(\omega^{\delta})$ and collapsing algebras
It is consistent that ${\mathbb{P}}(\omega^{\kappa+1})$ collapses $2^{\kappa}$ to $\omega_{1}$, although $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\kappa})<(2^{\kappa})^{+}$ and ${\mathbb{P}}(\omega^{\kappa})$ does not collapse $2^{\kappa}$ to $\omega$. First we note that the equality ${\mathfrak{h}}=\omega_{1}$ holds in Cohen extensions of models of GCH (see [3]). If $V^{\prime}$ is a Cohen model for ${\mathfrak{c}}=\omega_{5}$ obtained from a model $V$ of GCH, then in $V^{\prime}$ we have ${\mathfrak{h}}=\omega_{1}$ and, by Fact 2.4, $2^{\omega_{1}}=(\max(\omega_{5},2)^{\omega_{1}})^{V}=(\omega_{5}^{\omega_{1}})^{V}=\omega_{5}$. By Theorem 4.10(A) and Fact 2.6(b) we have $1_{{\mathbb{P}}(\omega^{\omega_{1}+1})}\Vdash|2^{\omega_{1}}|=\omega_{1}$ although (see [2]) $\mathop{\rm cc}\nolimits({\mathcal{P}}_{\omega_{1}})=\omega_{3}<\omega_{6}=(2^{\omega_{1}})^{+}$ and forcing by ${\mathbb{P}}(\omega^{\omega_{1}})$ does not collapse $2^{\omega_{1}}$ to $\omega$.
Theorem 5.9
$\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\mu})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$, if $\mu$ is a cardinal, $\mu>\mathop{\rm cf}\nolimits(\mu)=\omega$ and some of the following conditions hold
(a) $2^{\mu}=\mu^{+}$,
(b) $2^{<\mu}=\mu$,
(c) $\mathrm{MA}+\mu<{\mathfrak{c}}$,
(d) $V$ is a Cohen model for ${\mathfrak{c}}=\kappa$ obtained from a model $V^{\prime}$ of GCH in which $\kappa$ is regular and $\mu<\kappa$.
Then $\mathop{\rm ro}\nolimits(\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\mu+n})))\cong\mathop{\rm Col}\nolimits(\omega_{1},2^{\mu})$, for all $n\in{\mathbb{N}}$.
Proof. By Fact 2.9 we have $\omega^{\mu}=\mu$; thus ${\mathbb{P}}(\omega^{\mu})={\mathbb{P}}(\mu)=[\mu]^{\mu}$ and by Fact 2.1(d) we have $\mathop{\rm sq}\nolimits{\mathbb{P}}(\omega^{\mu})={\mathcal{P}}_{\mu}$. By Theorem 4.10 the poset ${\mathcal{P}}_{\mu}$ is $\omega_{1}$-closed and by Fact 2.6(c) collapses $\mu^{\omega}$ to $\omega_{1}$. So, by Fact 5.1 it is sufficient to prove that each of conditions implies that $\mu^{\omega}=2^{\mu}$.
If $2^{\mu}=\mu^{+}$, then, since by König’s theorem $\mu^{+}\leq\mu^{\omega}\leq 2^{\mu}$, we have $\mu^{\omega}=2^{\mu}$. If $2^{<\mu}=\mu$, then, since for a singular cardinal $\mu$ we have $2^{\mu}\in\{2^{<\mu}\mu^{\mathop{\rm cf}\nolimits(\mu)},(2^{<\mu})^{\mathop{\rm cf}\nolimits(2^{<\mu})}\}$, we obtain $\mu^{\omega}=2^{\mu}$. If $V\models\mathrm{MA}+\mu<{\mathfrak{c}}$, then ${\mathfrak{c}}\leq\mu^{\omega}\leq 2^{\mu}={\mathfrak{c}}$ and we have $\mu^{\omega}=2^{\mu}$. Finally, if $V=V^{\prime}_{\mathop{\mathrm{Fn}}\nolimits(\kappa,2)}[G]$, then by Fact 2.4 in $V$ we have $2^{\mu}=(\max(\kappa,2)^{\mu})^{V^{\prime}}=(\kappa^{\mu})^{V^{\prime}}=\kappa$ and $\mu^{\omega}=(\max(\kappa,\mu)^{\omega})^{V^{\prime}}=(\kappa^{\omega})^{V^{\prime}}=\kappa$; so, $\mu^{\omega}=2^{\mu}$ again. The second statement follows from the first and Theorem 5.7(b). $\Box$
Acknowledgement